OpenStudy (anonymous):

Breanna works in the mall and her hours increase during busy shopping times of the year. Breanna has a plan to save $25 each week for 5 weeks, and then increase that amount by $20 for the next 7 weeks during the busy time, then go back to $25 per week. If Breanna’s plan is to save $540, how many weeks total will she need to save? a. 11 weeks b. 16 weeks c. 22 weeks d. 23 weeks

OpenStudy (anonymous):

Find the total saved for the first 5 weeks. Then find the total saved for the next 7 weeks. Add the two amounts. Subtract that from the total desired....

OpenStudy (ranga):

Savings in the first 5 weeks: $25 x 5 = $125 Savings in the next 7 weeks: $45 x 7 = $315 Savings in the next "x" weeks: $25 x "x" = $25x Total savings: $125 + $315 + $25x = $540 440 + 25x = 540 25x = 540 - 440 = 100 x = 100/25 = 4 Total weeks = 5 + 7 + x = 12 + 4 = 16 weeks.